An organism of a certain type can grow from 30 to 195 in 5 hours. Find k for the growth formula.
Answer by stanbon(48558)

An organism of a certain type can grow from 30 to 195 in 5 hours. Find k for the growth formula.
---------------------------
You have two points: (0,30) and (5,195)
----------------------------
The model form is y = ab^x
Substitute to get:
195 = ab^5
30 = ab^0
--------------
Divide the 2nd equation into the 1st:
(195/3) = b^5
b = 2.3045...
-------------
From the 2nd equation, a = 30
----------------
EQUATION:
y = 30*2.3045^x
----------------
I'm not sure you mean by the "k" variable.
====================
Cheers,
Stan H.